OpenStudy (elainagalvan10.):

Graph the functions below. Arrange the parabolas with respect to the position of their vertices from left to right. y=f(4x+3) y=f(x-3) y=f(2x-1) y=f(x/2-2) y=f(x-7) y=f(x+3)/3 y=f(x+2) any help please!

OpenStudy (mathmale):

Suggestion: Assume that f(x) represents a parabola with vertex at (0,0). Draw it; assume that the function is f(x)=x^2. Now, determine how the graph of, say, y=f(x-3), would be different from that of y=f(x). In this case, the input x-3 to f(x) has a graph that is the same as that of f(x)=x^2, EXCEPT that it has been shifted 3 units to the right. If you have reference material available, look up "stretching and compression in graphing."